OpenStudy (anonymous):

radical (3x+y)=3+x^2y^2 find y' using implicit differentiation formula

OpenStudy (freckles):

Which part do you need help differentiating?

OpenStudy (freckles):

\[\frac{d}{dx}(x^2y^2) \text{ Note: Requires product rule }\]

OpenStudy (freckles):

oh it is \[\sqrt{3x+y}=3+x^2y^2 \\ \text{ Honestly I don't like \square roots so I'm going \to \square both sides } \\ 3x+y=(3+x^2y^2)^2\]

OpenStudy (anonymous):

1/2 radical(3x+y)(4+y') =2x^2.y.y' +2xy^2

OpenStudy (freckles):

But we could also do it your way...

OpenStudy (freckles):

except we need to use chain rule on the left hand side

OpenStudy (anonymous):

yeah, it's very hard to write here

OpenStudy (freckles):

\[\frac{1}{2 \sqrt{3x+y}} \cdot (3+y')=2xy^2 +2yy'x^2 \] \[\frac{3}{2 \sqrt{3x+y}}+\frac{y'}{2 \sqrt{3x+y}}=2xy^2 +2yy'x^2 \]

OpenStudy (freckles):

next step is to gather your y' parts together

OpenStudy (anonymous):

i don't know what you did on the right side

OpenStudy (freckles):

distribute?

OpenStudy (freckles):

a(b+c)=ab+ac

OpenStudy (freckles):

or (b+c)a=ab+ac

OpenStudy (freckles):